OpenStudy (anonymous):

the dimensions of a rectangle are so that the length is 9 in more than its width. if the length was doubled and the width was decreased by 4 in, the area would be increased by 110 in ^2. What are the length and the width of the rectangle?

OpenStudy (gorv):

let width=x

OpenStudy (gorv):

length=x+9 area=x(x+9)

OpenStudy (anonymous):

Area = Length x Width First rectangle A = 9 x. x Second rectangle A + 110 = (2x(9 x)) . (x-4 )

OpenStudy (gorv):

length is doubled=2*(x+9) width after decrement=x-4 new area=2*(x+9)*(x-4) new area - old area=110 now solve it